OpenStudy (anonymous):

subtracting fractions 6^3/8-5^1/4

OpenStudy (anonymous):

\[\frac{ 6^{3} }{8} - \frac{ 5^{1} }{ 4 }\] ?

OpenStudy (anonymous):

it is 6 and 3/8-5 and 1/4

OpenStudy (anonymous):

\[6\frac{ 3 }{8-5} - \frac{ 1 }{ 4 }\] ?

OpenStudy (anonymous):

6 and the fracion is 3/8 - 5 and fraction 1/4

OpenStudy (anonymous):

\[6 - \frac{ 3 }{8-5} - \frac{ 1 }{4}\] Lol, is this it?

OpenStudy (anonymous):

no the 5 goes with the 1/4

OpenStudy (anonymous):

\[6 - \frac{ 3 }{ 8} - 5\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

no minus between 6 and 3/8 then it will be correct

OpenStudy (anonymous):

six and three eights subtract five and one fourth ?

OpenStudy (anonymous):

ya and can you put the answer in simplist terms please

zepdrix (zepdrix):

\[\large 6\frac{3}{8}-5\frac{1}{4}\]

OpenStudy (anonymous):

\[\frac{ 204 }{ 32 } - \frac{ 168 }{ 32 } = \frac{ 36 }{ 32 } = \frac{ 9 }{ 8 }\]